# Outlier Gradient Analysis: Efficiently Improving Deep Learning Model Performance via Hessian-Free Influence Functions

### v2

@@ 4 Correctness Verification on Synthetic Data
Normalized Margin Northcutt et al. (2021) 82.0 89.0
Self-Confidence Müller and Markert (2019) 82.0 89.0
Confidence Entropy Kuan and Mueller (2022) 82.0 89.0
Exact Hessian Cook and Weisberg (1982) 90.0 90.0
Gradient Tracing Pruthi et al. (2020) 82.0 91.0
LiSSA Koh and Liang (2017) 82.0 91.0
DataInf Kwon et al. (2024) 82.0 91.0
Self-LiSSA Bejan et al. (2023) 82.0 90.0
@@
Outlier Gradient (iForest) 96.0 96.0
@@
We also conduct a quantitative evaluation to assess the advantages of our approach compared to three recently proposed noisy label correction methods and six influence function-based approaches, as detailed in Table 1. Specifically, we measure ground-truth outlier predictive accuracy and the performance gain achieved by removing detrimental samples. For noisy label correction approaches we consider: Normalized Margin (Northcutt et al., 2021), Self-Confidence (Müller and Markert, 2019), and Confidence-Weighted Entropy (Kuan and Mueller, 2022). The influence function approaches include computing the Hessian exactly (Cook and Weisberg, 1982), using the Hessian-free gradient tracing approach by (Pruthi et al., 2020), LiSSA-based optimization (Koh and Liang, 2017), the recently proposed influence estimation approach DataInf (Kwon et al., 2024), self-influence using LiSSA as in (Bejan et al., 2023), and self-influence using DataInf. We compute influences only using the training samples and performance is measured on the test set.
@@
Our outlier gradient analysis approaches demonstrate high accuracy in identifying mislabeled outliers (96-98%), outperforming all three noisy label correction baselines (only 82% accuracy) and among influence baselines, all exhibit similar performance except for exact Hessian computation, which attains 90% accuracy. Next, we evaluate model performance gain by removing detected outlier samples and retraining the MLP on the trimmed dataset. Here the benefits of our iForest outlier gradient analysis can be observed, as it increases performance from 90% to 96% while the overtly simple L1/L2-norm outlier analysis approaches are not as effective. The other baselines exhibit performance variations between 89-91%. This emphasizes the effectiveness of our iForest approach, while exhibiting low time complexity (refer to Appendix B.3 for details on computational complexity).
@@ 5 Noisy Label Correction for Vision Datasets
Normalized Margin Northcutt et al. (2021) 91.33 90.06 83.57 60.94
Self-Confidence Müller and Markert (2019) 91.38 90.09 83.65 60.51
Confidence Entropy Kuan and Mueller (2022) 91.11 90.05 83.63 60.62
Gradient Tracing Pruthi et al. (2020) 91.47 89.98 83.38 60.73
LiSSA Koh and Liang (2017) 91.49 90.05 83.38 60.48
DataInf Kwon et al. (2024) 91.46 90.05 83.40 60.70
Self-LiSSA Bejan et al. (2023) 92.07 89.58 83.01 59.48
@@
Here we demonstrate the effectiveness of our approach in addressing noisy label correction using the CIFAR-10N and CIFAR-100N real-world noisy label datasets (Wei et al., 2022). These datasets stem from the original CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009), but introduce label inaccuracies due to crowdsourced labeling. CIFAR-10N has 3 different noise settings: Aggregate, Random, and Worst– these correspond to obtaining the label using majority voting across 3 annotators, the first annotator label, and selecting the worst annotator label, respectively. CIFAR-100N only has a single noise setting.
Figure 2: Examples of detrimental noisy samples detected using our proposed outlier gradient analysis. Top row: CIFAR-10N; bottom row: CIFAR-100N. Top label (red): noisy label; bottom label (green): correct class.
@@
Table 2 shows the accuracy performance of our outlier gradient trimming approaches (iForest, L1/L2-norm) compared to label correction approaches and influence-based baselines covered in the previous section. Exact Hessian computation is excluded due to its computationally intractability for large datasets. Our outlier gradient analysis methods consistently outperform other baselines across diverse noise settings and datasets. Notably, even in challenging scenarios like the Worst noise setting in CIFAR-10N (40.21% noise rate), our approaches are the top performers– L1-norm based outlier analysis achieves highest accuracy gain, improving from 82.27% (vanilla ResNet-34) to 84.20%. Similar superior performance is observed in the Random noise setting (17.23% noise rate), where L2-norm outlier analysis achieves a final accuracy of 90.25% compared to original cross-entropy accuracy of 89.17% and in CIFAR-100N, where it attains the highest performance of 61.40%, surpassing the cross-entropy performance of 57.36%. In the CIFAR-10N Aggregate noise setting (noise rate 9.03%), outlier gradient analysis is again the top performer. Due to space constraints, we omit standard deviations from Table 2, but these are provided in Appendix B.1.
@@
Additionally, visual examples of mislabeled samples detected by our outlier gradient analysis approach (iForest) are provided in Figure 2. All displayed images contain mislabeled samples, and their removal from the training set contributes to improved model performance on the test set. In Table 2, we set the trimming budget for outlier gradient analysis ($k$) at 5% of the training data size. An empirical justification for this choice is provided in Appendix B.2, where we vary the outlier budget and measure test set accuracy. We also conduct ablations on the iForest parameters in Appendix B.4 and provide running time experiments in Appendix B.3 to benchmark the computational efficiency of our approach. We provide experiments with ResNet-18 He et al. (2016) as the base model in Appendix B.6 and on ImageNet Deng et al. (2009) in Appendix B.5 showing similar trends.
@@ 6 Data Selection for Fine-tuning NLP Models
We conduct experiments on data selection for fine-tuning on NLP models, following the experiment setup by Kwon et al. (2024) for DataInf, where the RoBERTa transformer model (Liu et al., 2019) is fine-tuned on four binary GLUE datasets (Wang et al., 2018): QNLI, SST2, QQP, and MRPC. More details on these datasets and RoBERTa are provided in Appendix A. To assess if influence-based methods can enhance NLP model performance via Low Rank Adaptation (LoRA) (Hu et al., 2022) fine-tuning, Kwon et al. (2024) introduce noisy versions of all four datasets by flipping the binary label for 20% randomly chosen training data samples.
The goal of the data selection task is to select the best representative subset of the training data so that performance is maximized on an unseen test set. Specifically, 70% of the most beneficial samples are selected according to each influence computation approach, and the model is fine-tuned for 10 epochs and rank of LoRA matrix is set to 4. Then, as the model trains over each epoch, performance is measured on the unseen test set. Clearly, for fairness, the sample influence is computed only using the training set, and the test set remains unknown until inference.
The results over three runs are presented in Figure 3 for all four GLUE datasets. We only show trends for iForest based outlier gradient analysis to aid visualization since performance is similar for the L1/L2-norm methods. It can be seen that our outlier gradient trimming approach markedly outperforms all other baselines but the Self-LiSSA (Bejan et al., 2023) self-influence baseline is competitive with our approach. More specifically, outlier gradient analysis achieves slightly better test set results on QNLI, SST2, QQP, and on MRPC, Self-LiSSA and outlier gradient analysis are on par with each other. Here, we would like to emphasize that despite competitive performance, our outlier gradient analysis is orders of magnitude faster than Self-LiSSA, as shown in experiments of Appendix B.3. These results highlight the effectiveness of our proposed outlier gradient analysis/trimming approach in selecting relevant data for fine-tuning NLP models while being more computationally efficient.
@@
Figure 4: Quantitative and qualitative results for outlier gradient analysis on LLM influential data identification benchmarks.
7 Influential Data Identification for LLMs
We demonstrate the effectiveness of our proposed outlier gradient analysis in identifying beneficial/detrimental samples for Large Language Models (LLMs), using benchmarks from DataInf (Kwon et al., 2024). Since LLMs are primarily employed for text generation tasks, and not text classification, we cannot directly measure the benefits of influence estimation as we did previously for RoBERTa using the GLUE datasets. Instead, Kwon et al. (2024) resort to an indirect measurement of influence for LoRA fine-tuned LLMs. Specifically, they seek to assess what training set prompts (used for fine-tuning) are identified as most influential for a given unseen test prompt. The robustness and effectiveness of influence estimation are gauged based on whether the identified training set prompts belong to the same class category as the given test prompt. We utilize the three benchmark datasets introduced in DataInf (Kwon et al., 2024): Sentence Transformations, Math Without Reasoning, and Math With Reasoning, to conduct the influential data identification experiment on the Llama-2-13B-chat [^2] https://ai.meta.com/llama/. LLM. For each of the influence identification benchmark datasets, there are 900 training samples for LoRA fine-tuning, and 10 categories or classes of task types with 90 samples belonging to each class. For each dataset there are 100 test set prompts and with 10 test set prompts per class category.
Table 3: AUC/Recall for outlier gradient analysis and baselines for influential class detection on 3 text generation tasks on Llama2-13B-chat LLM.
@@ 7 Influential Data Identification for LLMs
Sentence Transformations Gradient Tracing (Pruthi et al., 2020) 0.999 ± 0.001 0.982 ± 0.032
DataInf (Kwon et al., 2024) 1.000 ± 0.000 0.996 ± 0.012
Outlier Gradient Analysis 1.000 ± 0.000 1.000 ± 0.000
Math Problems Without Reasoning Gradient Tracing (Pruthi et al., 2020) 0.724 ± 0.192 0.241 ± 0.385
DataInf (Kwon et al., 2024) 0.999 ± 0.005 0.993 ± 0.046
Outlier Gradient Analysis 1.000 ± 0.000 1.000 ± 0.000
Math Problems With Reasoning Gradient Tracing (Pruthi et al., 2020) 0.722 ± 0.192 0.226 ± 0.376
DataInf (Kwon et al., 2024) 0.999 ± 0.004 0.990 ± 0.049
Outlier Gradient Analysis 1.000 ± 0.000 1.000 ± 0.000
@@
In Kwon et al. (2024), to predict the most influential training samples given a test set prompt, the authors assign a pseudo label to every data point in the training set (1 if it is in the same class/task category as the test data prompt, or 0 otherwise). This set serves as a ground-truth for measuring performance of influence functions in identifying influential data samples. Next, they calculate the Area Under the Curve (AUC) by comparing the absolute values of the influence function (for each training set prompt corresponding to a given test prompt) with these pseudo labels. Clearly, a high AUC signifies that training data samples from the same category have a significant influence on the given test prompt. The average AUC across all test data points is then recorded, and is denoted as the Class Detection (AUC) metric. Additionally, another metric is used– for every test data prompt, the authors determine if the proportion of training data prompts belonging to the same class/category are within the top 90 (# of training prompts in each category) influential samples. The average % across all test data points is calculated and this metric is denoted as Class Detection (Recall). A higher recall indicates more effective influence estimation.
@@
In our proposed outlier gradient analysis approach, we initially used outlyingness labels directly for identifying detrimental samples. However, for a more fine-grained analysis, we now require detailed outlyingness values. To achieve this, we require outlyingness scores and hence, train 10 individual iForest estimators for each class prompt category, as the ultimate objective is to use outlier gradient analysis for prompt class detection. Each class’s iForest estimator is trained solely on the gradient space of training prompts from that category. Subsequently, for each test set prompt, we utilize each iForest estimator to generate an outlier score based on the gradient space of that test sample. This enables us to conduct the influential data identification experiment for our proposed method.
@@
The results of this experiment are presented in Table 3. Our outlier gradient analysis performs exceptionally well on this task, achieving perfect scores for both AUC and Recall. It outperforms DataInf and Gradient Tracing, with LiSSA omitted as it fails to converge due to instability on LLMs (Kwon et al., 2024). Self-influence baselines also cannot be used since a similarity matrix with test prompts needs to be constructed. Figure 4 further illustrates the individual influence predictions, with darker colors indicating lower outlier score magnitudes. The heatmaps correspond to each of the three benchmark datasets, with test samples ordered sequentially based on their class category. The accurate influence estimation is evident from the highest influence values along the diagonal. The most influential sample identified by our approach closely resembles the given test prompt for each of the three benchmark datasets, showcasing the benefits of employing outlier gradient analysis in LLMs.

### v3

@@ 4 Hypothesis Verification on Synthetic Data
Normalized Margin 82.0 89.0
Self-Confidence 82.0 89.0
Confidence Entropy 82.0 89.0
Exact Hessian 90.0 90.0
Gradient Tracing 82.0 91.0
LiSSA 82.0 91.0
DataInf 82.0 91.0
Self-LiSSA 82.0 90.0
@@
We seek to validate the hypothesis of our proposed idea and showcase the effectiveness of our outlier gradient analysis method on two synthetic 2D toy datasets [^1] Comprehensive details regarding datasets and model training for experiments are provided in Appendix B. and two models for binary classification in Figure 1. In this figure, subfigures A-D present a linear dataset employing a Logistic Regression model, while subfigures E-H exhibit a non-linear dataset utilizing a non-convex Multilayer Perceptron (MLP) model as the base model. Specifically, subfigures A and B depict the training and test sets of a linearly separable dataset comprising 150 and 100 samples, respectively. Notably, the training set includes 10 manually generated noisy samples with misspecified labels. Subfigure C displays the influence score of each training sample, computed using Eq. (1), and subfigure D provides a visualization of the gradient space. Similarly, subfigures E and F represent the training and test sets of the two half moons dataset, with the training set consisting of 250 samples and the test set of 100 samples, equally distributed between two classes. The training set in this case also contains 20 noisy samples. Subfigures G and H showcase the influence score and gradient space of the non-convex case.
Figure 1: Illustrating our outlier gradient analysis approach on two synthetic datasets and convex/non-convex models. A-D showcase our outlier gradient analysis approach on a 2D linearly separable synthetic dataset. This dataset includes a small subset of detrimental samples with incorrect labels used to train a Logistic Regression binary classification model. Meanwhile, E-H depict our outlier gradient analysis on a non-linear synthetic dataset with mislabeled samples employed in training a Multilayer Perceptron (MLP) neural network. In subfigures A and E, the training sets are represented with class labels 0 (red) and 1 (blue) in the convex and non-convex cases, respectively. Detrimental samples with incorrect class labels are marked with $\times$, while regular samples are marked with $\circ$. B and F denote the test sets used to evaluate model performance. C and G display the influence scores calculated by Eq. (1). Note that G demonstrates that influence scores are not reliable indicators for detecting detrimental samples in the non-convex case. After applying outlier analysis on the gradient space of the non-convex MLP model, most detrimental samples are detected. D and H showcase the gradient space obtained for each sample from the Logistic Regression and MLP models, respectively. It is evident that the outlier samples correspond to detrimental samples with mislabeled classes, which are linearly or non-linearly separated from inliers. Note that the benefits of outlier gradient trimming can be clearly observed—removing predicted outlier samples via iForest and retraining the MLP enhances classification performance from $90\%\rightarrow 96\%$ on the test set (refer to Table 1).
In the linear case, as illustrated in subfigure C, the influence score proves to be a reliable indicator for distinguishing detrimental samples from beneficial ones. Notably, detrimental samples exhibit large negative scores, while other samples display positive or nearly zero values. Additionally, subfigure D affirms that these detrimental samples are distinctly separated in the gradient space, confirming the validity of the equivalent transformation outlined in Hypothesis 3.2. However, the limitations of influence scores become evident in the context of non-convex models, as observed in subfigure G, where the influence scores of detrimental samples are mixed with those of normal ones. Nevertheless, in the gradient space illustrated in subfigure H, the detrimental samples are effectively isolated from inliers. Notably, our method does not rely on the Hessian for computing influence and operates directly on the gradient space using outlier analysis.
@@
We also conduct a quantitative evaluation to assess the advantages of our approach compared to three recently proposed noisy label correction methods and six influence function-based approaches, as detailed in Table 1. Specifically, we measure ground-truth outlier predictive accuracy and the performance gain achieved by removing detrimental samples. For noisy label correction approaches we consider: Normalized Margin (Northcutt et al., 2021), Self-Confidence (Müller and Markert, 2019), and Confidence-Weighted Entropy (Kuan and Mueller, 2022). The influence function approaches include computing the Hessian exactly (Cook and Weisberg, 1982), using the Hessian-free gradient tracing approach [^2] We only use the last checkpoint in Gradient Tracing (Pruthi et al., 2020) for fair comparisons. by (Pruthi et al., 2020), LiSSA-based optimization (Koh and Liang, 2017), the recently proposed influence estimation approach DataInf (Kwon et al., 2024), self-influence using LiSSA as in (Bejan et al., 2023), and self-influence using DataInf. We compute influences only using the training samples and performance is measured on the test set.
@@
Our outlier gradient analysis approaches demonstrate high accuracy in identifying mislabeled outliers (96-98%), outperforming all three noisy label correction baselines (only 82% accuracy) and among influence baselines, all exhibit similar performance except for exact Hessian computation, which attains 90% accuracy. Next, we evaluate model performance gain by removing detected outlier samples and retraining the MLP on the trimmed dataset. Here the benefits of our iForest outlier gradient analysis can be observed, as it increases performance from 90% to 96% while the overtly simple L1/L2-norm outlier analysis approaches are not as effective. The other baselines exhibit performance variations between 89-91%. This emphasizes the effectiveness of our iForest approach, while exhibiting low time complexity (refer to Appendix C.3 for details on computational complexity).
@@ 5 Noisy Label Correction for Vision Datasets
Here we demonstrate the effectiveness of our approach in addressing noisy label correction using the CIFAR-10N and CIFAR-100N real-world noisy label datasets (Wei et al., 2022). These datasets stem from the original CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009), but introduce label inaccuracies due to crowdsourced labeling. CIFAR-10N has 3 different noise settings: Aggregate, Random, and Worst– these correspond to majority voting across 3 annotators, the first annotator label, and selecting the worst annotator label, respectively. CIFAR-100N only has a single noise setting.
@@
Normalized Margin 91.33 90.06 83.57 60.94
Self-Confidence 91.38 90.09 83.65 60.51
Confidence Entropy 91.11 90.05 83.63 60.62
Gradient Tracing 91.47 89.98 83.38 60.73
LiSSA 91.49 90.05 83.38 60.48
DataInf 91.46 90.05 83.40 60.70
Self-LiSSA 92.07 89.58 83.01 59.48
@@
Table 2 shows the accuracy performance of outlier gradient analysis (L1/L2-norm, iForest) compared to label correction approaches and influence-based baselines covered in the previous section. Exact Hessian computation is excluded due to its computational intractability for large datasets. Our outlier gradient analysis methods consistently outperform other baselines across diverse noise settings and datasets. Notably, even in challenging scenarios like the Worst noise setting in CIFAR-10N (40.21% noise rate), our approaches are the top performers– L1-norm based outlier analysis achieves highest accuracy gain, improving from 82.27% (vanilla ResNet-34) to 84.20%. Similar superior performance is observed in the Random noise setting (17.23% noise rate), where L2-norm outlier analysis achieves a final accuracy of 90.25% compared to original cross-entropy accuracy of 89.17% and in CIFAR-100N, where it attains the highest performance of 61.40%, surpassing the cross-entropy performance of 57.36%. In the CIFAR-10N Aggregate noise setting (noise rate 9.03%), outlier gradient analysis is again the top performer. Due to space constraints, we omit standard deviations from Table 2, but these are provided in Appendix C.1.
@@
Additionally, visual examples of mislabeled samples detected by our outlier gradient analysis approach (iForest) are provided in Figure 2. All displayed images contain mislabeled samples, and their removal from the training set contributes to improved model performance on the test set. In Table 2, we set the trimming budget for outlier gradient analysis ($k$) at 5% of the training data size. An empirical analysis for the choice of $k$ is undertaken in Appendix C.2, where we vary the outlier budget (from 2.5% to 12.5%) and measure test set accuracy across the CIFAR-10N dataset.
Figure 2: Some detrimental samples detected using our proposed outlier gradient analysis. Top row: CIFAR-10N; bottom row: CIFAR-100N. Top label (red): noisy label; bottom label (green): correct class.
Additional Analyses. We conduct ablations on the iForest parameters in Appendix C.4. Further, we provide running time experiments on CIFAR-10N and CIFAR-100N in Appendix C.3 along with the other baselines. We also provide results with ResNet-18 as the base model in Appendix C.5 and on ImageNet (Deng et al., 2009) in Appendix C.6, showing similar trends. Finally, approaches for noisy learning can be categorized into methods that either change the loss function or model architecture or methods that identify noisy samples and remove/relabel them for improving performance (Algan and Ulusoy, 2021). Since our approach belongs to the latter category, we only compare against other approaches from this category. For completeness, we also present results comparing our approach with some others in the former category in Appendix C.7. We would like to emphasize that this is not an exhaustive list of baselines and noisy learning by adjusting the loss/model is not the focus of our work (but detecting detrimental samples is). Moreover, our algorithm could also be combined with approaches from both categories for additional gains.
@@ 6 Data Selection for Fine-tuning NLP Models
We conduct experiments on data selection for fine-tuning on NLP models, following the experimental setup by Kwon et al. (2024) for DataInf, where the RoBERTa transformer model (Liu et al., 2019) is fine-tuned on four binary GLUE datasets (Wang et al., 2018): QNLI, SST2, QQP, and MRPC. To assess if influence-based methods can enhance NLP model performance via Low Rank Adaptation (LoRA) (Hu et al., 2022) fine-tuning, Kwon et al. (2024) introduce noisy versions of all four datasets by flipping the binary label for 20% randomly chosen training data samples. The goal of the data selection task is to select the best representative subset of the training data so that performance is maximized on an unseen test set. Specifically, 70% of the most beneficial samples are selected according to each influence computation approach, and the model is fine-tuned for 10 epochs and rank of LoRA matrix is set to 4. Then, as the model trains over each epoch, performance is measured on the unseen test set. Clearly, for fairness, the sample influence is computed only using the training set, and the test set remains unknown until inference.
@@
The results over three runs are presented in Figure 3 for all four GLUE datasets. We only show trends for iForest based outlier gradient analysis to aid visualization since performance is similar for the L1/L2-norm methods. It can be seen that our outlier gradient trimming approach markedly outperforms all other baselines but the Self-LiSSA (Bejan et al., 2023) self-influence baseline is competitive with our approach. More specifically, outlier gradient analysis achieves slightly better test set results on QNLI, SST2, QQP, and on MRPC, Self-LiSSA and outlier gradient analysis are on par with each other. Here, we would like to emphasize that despite competitive performance, our outlier gradient analysis is orders of magnitude faster than Self-LiSSA, as shown in experiments of Appendix C.3. These results highlight the effectiveness of our proposed outlier gradient analysis/trimming approach in selecting relevant data for fine-tuning NLP models while being more computationally efficient.
Figure 4: Results for outlier gradient analysis on LLM influential data identification benchmarks.
7 Extending to Influential Data Identification for LLMs
We now consider an alternate task– demonstrating the effectiveness of our proposed outlier gradient analysis in identifying influential data samples for Large Language Models (LLMs), using the proposed benchmarks from DataInf (Kwon et al., 2024). The LLM influential data identification task at its core is a similarity measurement task, as it seeks to ascertain which fine-tuning prompts are most similar to a given test sample. More specifically, the goal is to assess what training set prompts (used for LoRA fine-tuning) are most influential for a given unseen test prompt. The robustness and effectiveness of influence estimation are gauged based on whether the identified training set prompts belong to the same class category as the given test prompt. We utilize the three benchmark datasets introduced in DataInf (Kwon et al., 2024): Sentence Transformations, Math Without Reasoning, and Math With Reasoning, to conduct the influential data identification experiment on the Llama-2-13B-chat [^3] https://ai.meta.com/llama/. LLM. For each of the influence identification benchmark datasets, there are 900 training samples for LoRA fine-tuning, and 10 categories or classes of task types with 90 samples belonging to each class. For each dataset there are 100 test set prompts with 10 test set prompts per class category.
Table 3: AUC/Recall for outlier gradient analysis and baselines for influential class detection on three text generation tasks on Llama2-13B-chat LLM.
@@ 7 Extending to Influential Data Identification for LLMs
Sentence Transformations Gradient Tracing 0.999 ± 0.001 0.982 ± 0.032
DataInf 1.000 ± 0.000 0.996 ± 0.012
Outlier Gradient 1.000 ± 0.000 1.000 ± 0.000
Math Problems Without Reasoning Gradient Tracing 0.724 ± 0.192 0.241 ± 0.385
DataInf 0.999 ± 0.005 0.993 ± 0.046
Outlier Gradient 1.000 ± 0.000 1.000 ± 0.000
Math Problems With Reasoning Gradient Tracing 0.722 ± 0.192 0.226 ± 0.376
DataInf 0.999 ± 0.004 0.990 ± 0.049
Outlier Gradient 1.000 ± 0.000 1.000 ± 0.000
@@
In Kwon et al. (2024), to predict the most influential training samples given a test set prompt, the authors assign a pseudo label to every data point in the training set (1 if it is in the same class/task category as the test data prompt, or 0 otherwise). This set serves as a ground-truth for measuring performance of identifying influential data samples. Next, they calculate the Area Under the Curve (AUC) by comparing the absolute values of the influence function (for each training set prompt corresponding to a given test prompt) with these pseudo labels. Clearly, a high AUC signifies that training data samples from the same category have a significant influence on the given test prompt. The average AUC across all test data points is then recorded, and is denoted as the Class Detection (AUC) metric. Additionally, another metric is used– for every test data prompt, the authors determine if the proportion of training data prompts belonging to the same class/category are within the top 90 (# of training prompts in each category) influential samples. The average % across all test data points is calculated and this metric is denoted as Class Detection (Recall), where higher recall is better.
@@
As part of this task, we need to measure similarity between train and test set samples. Note that for our experiments on identifying detrimental samples outlier gradient analysis only operated on the training set (i.e., it uses the training set gradients). However, to extend outlier analysis to this task while maintaining consistency with the previous experiments and methods, we will train 10 individual iForest estimators for each class prompt category, as the ultimate objective is to use outlier gradient analysis for prompt class detection. Each class’s iForest estimator is trained solely on the gradient space of training prompts from that category. Subsequently, for each test set prompt, we utilize each iForest estimator to generate an outlier score based on the gradient space of that test sample. This enables us to conduct the influential data identification experiment for our proposed method. The other baseline influence methods already have access to the given test set sample and can use that information directly for analyzing which training sample is most influential.
@@
Our outlier gradient analysis performs exceptionally well on this task, achieving perfect scores for both AUC and Recall in Table 3. It outperforms DataInf and Gradient Tracing, with LiSSA omitted as it fails to converge due to instability on LLMs (Kwon et al., 2024). Self-influence baselines also cannot be used since a similarity matrix with the full set of test prompts needs to be constructed (information leakage). Figure 4 further illustrates the individual influence predictions, with darker colors indicating lower outlier score magnitudes. The heatmaps correspond to three benchmark datasets, with test samples ordered sequentially based on their categories. The accurate influence estimation is evident from the highest influence values along the diagonal. The most influential sample identified by our approach closely resembles the given test prompts.
8 Discussion
Computational complexity and running time. Throughout, we have emphasized that outlier gradient analysis is efficient while being highly accurate at identifying detrimental training samples. We also conduct experiments to validate this empirically. In Table 6 (Appendix C.3), we benchmark the running time for all the methods considered for the various noise settings of CIFAR-10N and CIFAR-100N. It can be observed that outlier gradient analysis features in the top-performing methods in terms of computational efficiency, while simultaneously also featuring as a top-performing method for accurately detecting detrimental samples (as seen in Table 2). We observe similar trends for the ImageNet dataset in Table 10 (Appendix C.6). Note that this is also evident in terms of worst-case computational complexity, as outlier gradient analysis possesses linear (in both number of samples and parameters) time complexity (see Table 7 in Appendix C.3 for more details).
Adapting outlier gradient analysis to a validation/test set distribution. In some scenarios we might wish to utilize a validation set distribution to accurately adjust influence estimation. This is especially true for distribution shift scenarios, where the training and validation distributions are different. In the original influence formulation, the first term provides this information. For outlier gradient analysis, we only use training set gradients. To rectify this, we can instead employ a semi-supervised outlier analysis algorithm $\mathcal{A}$ with validation samples provided as inliers. We utilize the semi-supervised OneClassSVM (Li et al., 2003) outlier analysis algorithm and the distribution shift experimental framework from Chhabra et al. (2024) to assess performance. These results indicate that outlier gradient analysis is the top-performer across baselines, as can be seen in Table 12 (Appendix C.8). While a full extensive analysis of validation set adaptation is beyond the scope of this paper, these preliminary experiments showcase the benefits of outlier gradient analysis beyond just the training distribution.